Circular No. 7608

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE B. E. Schaefer, University of Texas at Austin, on behalf of the QUEST

collaboration (cf. IAUC 7387), reports on the discovery of twelve apparent supernovae discovered with the QUEST 16-CCD array camera at Llano del Hato. The total area searched was 254 deg^{2} between Mar. 15 and 25 to R = 20.2. Nightly scans continued until Apr. 4, so that the twelve objects all have well-sampled 3-color light curves over roughly three weeks, and many more supernovae are expected to be found. Each apparent supernova was found by subtraction of QUEST reference images taken between Mar. 3 and 8 and each new object has been confirmed via five or more independent images taken on three or more nights with subtractions from three independent reference images. Spectroscopy is still needed for SNe 2001am, 2001as, 2001at, 2001au, 2001av, 2001aw, and 2001ax for identifying supernova type and redshift; though spectroscopy has been obtained for the other objects, their types and redshifts are not yet known.

SN  2001 UT  α_{2000}  δ_{2000}  R  Offset
2001am Mar. 15  9^{h}11^{m}47.^{s}43 − 1^{o}18′58′′.2 19.1 0′′ E, 0′′ N
2001an Mar. 15 10 05 35.76 − 1 28 42.0 19.1 3′′ W, 1′′ S
2001ao Mar. 15 10 13 45.70 − 1 24 51.8 19.9 1′′ W, 1′′ S
2001ap Mar. 15 12 18 23.32 − 1 54 48.5 18.9 0′′ E, 0′′ N
2001aq Mar. 15 14 01 33.76 − 0 59 48.7 19.9 1′′ W, 0′′ N
2001ar Mar. 15 14 40 53.46 − 0 39 24.6 18.7 0′′ E, 1′′ S
2001as Mar. 16 15 09 45.32 − 0 55 38.0 19.1 1′′ W, 0′′ N
2001at Mar. 16 15 10 09.06 − 0 58 29.2 18.8 0′′ E, 0′′ N
2001au Mar. 16 15 16 49.17 − 0 07 46.6 18.6 3′′ E, 3′′ N
2001av Mar. 23 15 43 02.92 − 2 02 20.0 20.2 4′′ E, 0′′ N
2001aw Mar. 20 15 55 32.11 − 0 10 54.6 19.5 0′′ E, 0′′ N
2001ax Mar. 20 16 07 13.69 − 0 04 49.2 17.5 0′′ E, 7′′ S

Corrigendum. On IAUC 7391, line 1, for Further to IAUC 7391, read Further to IAUC 7387,

COMET 45P/HONDA-MRKOS-PAJDUŠÁKOVÁ On Apr. 4.42 UT, K. Kadota, Ageo, Saitama, Japan, reports m_{1} =

10.5 and coma diameter 0′.8 (0.18-m reflector + CCD). An ephemeris for 45P appears on MPC 42593 (orbital elements on MPC 42548).

2001 April 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green